Two bodies of masses M1 and M2 have equal energies.If P1 and P2 are their respective momentum,compute P1/P2.


1
Expert's answer
2021-05-07T17:27:38-0400

"E_1=E_2\implies\frac{m_1v_1^2}{2}=\frac{m_2v_2^2}{2}\implies \frac{v_1}{v_2}=\sqrt{\frac{m_2}{m_1}},"

"\frac{p_1}{p_2}=\frac{m_1v_1}{m_2v_2}=\frac{m_1}{m_2}\cdot \sqrt{\frac{m_2}{m_1}}=\sqrt{\frac{m_1}{m_2}}."